OpenStudy (anonymous):

let a & b be events with P(A) = 0.3 P(B) = 0.5 & P(anb)=0.2 probability that a or b happens a does not happen neither a or b happens either

OpenStudy (anonymous):

P(A or B) = P(A) + P(B) - P(A and B) P(not A) = 1 - P(A) P(neither A nor B) = 1 - P(A or B)